OpenStudy (vera_ewing):

Question 1: Titrations

OpenStudy (vera_ewing):

Copper metal is easily oxidized to Cu2+ by nitric acid according to the following equation: 3Cu(s) + 8HNO3(aq) → 3Cu(NO3)2(aq) + 2NO(g) + 4H2O(l) A copper penny with a mass of 3.067 g is dissolved in 100.0 mL of 1.000 M nitric acid. a. How many moles of copper are in the penny assuming it is pure copper? (5 points) b. When the reaction stops, the undissolved penny is removed. What is the mass of the undissolved penny? (6 points) c. The solution prepared in step B is transferred quantitatively to a 250.00 mL volumetric and diluted to the line. What is the molarity of the Cu2+ ion in this solution? (4 points)

OpenStudy (gebooors):

First calculate amount of substance of nitric acid

OpenStudy (gebooors):

I think you manageri a) easily :)

OpenStudy (vera_ewing):

so what would the answers for a b and c? i want to see if they line up with my answers

OpenStudy (vera_ewing):

just let me know what you get and i'll see if they match up with mine :)

OpenStudy (gebooors):

@Hoslos can U continue?

OpenStudy (anonymous):

You can also first find the moles of copper using the formula n=m/Mr, where n is moles, m is mass and Mr is given mass. So n= 3.067/64= 0.047 moles.

OpenStudy (anonymous):

Yes, I can.

OpenStudy (gebooors):

B) I wonder is there enough acid to dissolve all copper or Is acid limiting reactant?

OpenStudy (gebooors):

Thank you @Hoslos

OpenStudy (gebooors):

I see that there is not enough acid from equation and number @Hoslos got. Have to go now, bye for the next time and sorry

OpenStudy (anonymous):

Ok, for that, you have to back to the question and check the molar ration of Cu as to HNO3. It is 3:8, but for this exercise you have 0.047, which is as to x. Making x subject of the formula is x= 0.047*8/3= 0.125mol of HNO3. This is the molarity of the acid expected. But according to the question, there is undissolved penny, meaning the acid must be limiting, so basing from its given volume and concentration, let us find the moles contained. For liquids, moles=CV/1000, where C id concentration, V volume in cm^3. Hence n= 1*100/1000= 0.1mol. So this is the moarity of acid available. To know the reacted penny's amount, we need to find the molar basing from this available moles of the limiting acid, being 3:8 again and x:0.1 then x=0.1*3/8=0.037 mol of penny reacted fully. Thereore the unreacted moles are 0.047-0.037= 0.010mol.

OpenStudy (anonymous):

No problem @Gebooors .

OpenStudy (vera_ewing):

ok great. and for b) and c) ?

OpenStudy (anonymous):

What I explained now, Vera was b).

OpenStudy (anonymous):

Are you ok up to this point?

OpenStudy (vera_ewing):

And what about a) and c) ?

OpenStudy (anonymous):

I answered a), when Geboors first mentioned my name in this chat, can you see?

OpenStudy (vera_ewing):

oh yes i see. thanks. now just c)

OpenStudy (anonymous):

Ok. Let me carry on then.

OpenStudy (vera_ewing):

ok

OpenStudy (anonymous):

Well, I am not familir with those words, but from what I understood, when something is transfered quantitatively, it means that have made the ratio of moles as to the ration of volume. For instance, previously, the volume of acid was 100mL and then changed to 250mL, meaning it multiplied by 2.5, hence the moles of penny should also multiply in this ration, being 0.037 * 2.5 = 0.009 moles.

OpenStudy (vera_ewing):

so is that the answer for c) ?

OpenStudy (anonymous):

Yes.

OpenStudy (vera_ewing):

thanks so much! :)

OpenStudy (anonymous):

Again, it is My pleasure.

OpenStudy (gebooors):

@Hoslos, good work!

OpenStudy (anonymous):

Thanks @Gebooors .